OpenStudy (anonymous):

what is the minimum volume of a block of wood (density = 850 kg/m3) if it is to hold a 50 kg person entirely above the water when he stands on it?

OpenStudy (anonymous):

guys help its urgent

OpenStudy (anonymous):

block of 1m^3 would displace 850kg of water i addition a person stands on it its mass will be 900kg so it will displace 900kg of water. so in this case person would be perfectly above water. . . so formula should be mass of block+mass of man=900 so the block should displace atleast 900kg of water so its volume must be around or less than 0.9m^3

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

ok kanishk let me specify why i said less than 0.9m^3 because 0.9m^3 means 850*0.9= 765kg is the mass of block +50kg of mass of man tha is close to 815kg and water displaced is stil 900kgmuch more than required. . . so it must be less than 0.9m^3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

0.850*x+50=x 0.150=50 x=50*1000/150=1000/3 this is the volume of block minimum so that man can float on block. . .